An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 796 hours and a standard deviation of 55 hours. Find the probability that a random samples of 40 bulbs will have an average life between 788 and 810 hours.

1
Expert's answer
2022-02-02T10:09:56-0500

"\mu=796\\\sigma=55\\n=40"

We find the probability,

"p(788\lt \bar x\lt 810)"

We standardize this probability to a standard normal distribution as below.

"p(788\lt \bar x\lt 810)=p({788-\mu\over{\sigma\over \sqrt {n}}}\lt {\bar x-\mu\over{\sigma\over\sqrt{n}}}\lt {810-\mu\over{\sigma\over \sqrt {n}}})". Substituting for the values of "\mu,\sigma" and "n" given above, we have,

"p({788-796\over{55\over \sqrt {40}}}\lt Z\lt {810-796\over{55\over \sqrt {40}}})=P(-0.92\lt Z\lt 1.61)=\phi(1.61)-\phi(-0.92)=0.9463-0.1788=0.7675"

Therefore, the probability that a random sample of 40 bulbs will have an average life between 788 and 810 hours is 0.7675.